Some correspondences of principles in interpretability logic Mladen Vuković, Zagreb, Croatia

Keywords: interpretability logic, generalized Veltman semantics AMS Subject classification: 03F25 Abstract: A generalized Veltman semantics developed by de Jongh is used to investigate relations between several extensions of intepretability logic, IL.